OpenStudy (anonymous):

find the sum and product of the roots x^2+8x-30=0

OpenStudy (anonymous):

sum=-b/a product=c/a

OpenStudy (anonymous):

sum=-8/1=-8 product=-30/1=-30

OpenStudy (anonymous):

Thanks